1.Given a solution in which [Cu2+] = 0.20 mol/L and [NH3] = 1.40 mol/L, what is the equilibrium concentration of free Cu2+? The complex Cu(NH3)42+ has Kf = 5.0 X 1013. The answer to the question is 3.1 X 10 -14 I just need how to do it with the work shown.

2. You titrate a 20.00 mL sample of ammonia (Kb = 1.8 X 10-5) to the equivalence point using 10.00 mL of .20 M HCl solution. What is the pH of the solution at the equivalence point of the titration. The answer to the question is 5.21 but I do not know how to get there.
Answer #1

Bolution: Ô olution [Cu?] = 0.2 molle [Nm] = 1.40 mol/L 2 . Cu toa) + 4 NMD lens = [Cu (N12)4] : Ef = 5-oxiol Limiting ReageIcy (cy ²+] =...0.2 J = 0.648 X10S Trdi TTHC [C42+] = 1 Cu2+] = (C4) = 0.3086 X1013 3.08 xron!4 3.1 Xroly Answer M solutionph=7-12 log (60667) - 12 X 4.74 P) = 1 - x (-1-12t) - 237 (pn = 5.21) Angweni Scanned with CamScanner